# ECG-Byte: A Tokenizer for End-to-End Generative Electrocardiogram Language Modeling

### v1

@@ 3.2 ECG as Bytes
After training ECG-Byte, we encode any quantized ECG signal $X_{\text{symb}}$ by first converting each byte in the ECG signal to a 32-bit unsigned integer and building a trie structure, where each node represents a byte or a merged token sequence from prior encoding steps. The trie is initialized with single-byte tokens (0-255) and is extended with custom token sequences from the learned merge history. For each byte sequence in the input, the encoding function traverses the trie to find the longest match, replacing matched sequences with their assigned token IDs. If no match is found, the byte is added to the output as-is. The final encoded sequence is returned as output_ids, where we will denote as $X_{\text{ID}}$. We provide the detailed pseudocode of the encoding process in Algorithm 2.
Input: Input $X_{concat}$, Number of merges num_merges
Output: Tuple containing final encoded IDs, vocabulary map, and merge history
Function byte_pair_encoding($X_{concat}$, num_merges):
       Convert $X_{concat}$ to ID vector ids by casting each byte to $u32$;
       Initialize vocab with mappings from IDs 00 to $255$ to their string representations;
       Initialize vocab_tokens with mappings from IDs 00 to $255$ to singleton lists;
       Initialize empty list merges;
      for $i\leftarrow 0$ to num_merges exclusive do
             pairs $\leftarrow$ get_stats(ids) using parallel processing;
            if pairs is empty then
                   break
             end if
            best_pair $\leftarrow$ Pair in pairs with highest frequency;
             if best_pair is not found then
                   break
             end if
            new_id $\leftarrow 256+i$;
             ids $\leftarrow$ merge(ids, best_pair, new_id);
            vocab[new_id] $\leftarrow$ concat(vocab[best_pair.0], vocab[best_pair.1]);
            new_token $\leftarrow$ concat(vocab_tokens[best_pair.0], vocab_tokens[best_pair.1]);
             vocab_tokens[new_id] $\leftarrow$ new_token;
            merges.append(new_token, new_id);
       end for
      return (ids, vocab, merges);
Algorithm 1 Training Process for ECG-Byte
Input: Input $X_{\text{symb}}$, Merge history merges containing pairs of token sequences and their token IDs
Output: Vector of encoded IDs
Function encode($X_{\text{symb}}$, merges):
       ids $\leftarrow$ Convert $X_{\text{symb}}$ to vector of u32 by casting each byte;
       Initialize root TrieNode trie_root using TrieNode::new();
      for b $\leftarrow 0$ to $255$ do
             insert(trie_root, [b], b);
       end for
      foreach (token_sequence, token_id) in merges do
             insert(trie_root, token_sequence, token_id);
       end foreach
      Initialize empty list output_ids;
       i $\leftarrow 0$;
      while i $<$ length of ids do
             node $\leftarrow$ trie_root;
             match_len $\leftarrow 0$;
             match_id $\leftarrow$ None;
            for j $\leftarrow$ i to length of ids do
                   id $\leftarrow$ ids[j];
                  if id exists in node.children then
                         node $\leftarrow$ node.children[id];
                        if node.token_id is not None then
                               match_len $\leftarrow$ j - i + 1;
                               match_id $\leftarrow$ node.token_id;
                         end if
                   end if
                  else
                         break
                   end if
             end for
            if match_id is not None then
                   output_ids.append(match_id);
                   i $\leftarrow$ i + match_len;
             end if
            else
                   output_ids.append(ids[i]);
                   i $\leftarrow$ i + 1;
             end if
       end while
      return output_ids;
Algorithm 2 Encoding Process for ECG-Byte
@@ 3.3 Large Language Model
In this study, we utilize the Llama-3.2-1B (Grattafiori et al., 2024) checkpoint through the HuggingFace API (Wolf et al., 2020) unless specified otherwise. The Llama 3.2 series model is a variant of the Llama 3 models (Grattafiori et al., 2024) and support context lengths of up to 128k tokens. They are notable for their superior performance despite having 1 billion parameters, making them highly efficient and capable models to test our methodology upon. We also provide an ablation study in subsection 5.4 where we utilize other popular LLMs, such as GPT2 XL 1.5B (Radford et al., 2019), Gemma 2B (Team et al., 2024), and OPT 1.3B (Zhang et al., 2022).
@@ 3.4 Learning Objective
The learning objective for training the LLM considers a sequence composed of three parts, $\{X_{\text{ID}},Q,\mathcal{S}\}$, where $X_{\text{ID}}\in\mathcal{V}^{M}$ represents the encoded ECG sequence of length $l_{X_{\text{ID}}}=|X_{\text{ID}}|$, with each token drawn from the extended vocabulary $\mathcal{V}$ of size $M$, $Q$ represents the tokenized question, and $\mathcal{S}$ denotes the tokenized answer sequence. The input sequence includes special tokens: - [BOS] as the beginning-of-sequence token, - [SIG_START] and [SIG_END] to indicate the start and end of the encoded ECG sequence, and - [EOS] as the end-of-sequence token for the generated answer. The motivation for adding [SIG_START] and [SIG_END] special tokens is inspired by Liu et al. (2023), where they utilize special tokens indicating the start and end of the image. Thus, the full input sequence is structured as:
[…]
$L=1+1+l_{X_{\text{ID}}}+1+l_{Q}+l_{\mathcal{S}}+1,$
@@ 4.1 Experimental Settings
We fine-tuned the LLM using the AdamW optimizer (Kingma and Ba, 2017) with a learning rate of $41e-$ 4, weight decay of $21e-$ 2, and a custom learning rate scheduler. This scheduler applies an initial learning rate $\mathrm{init\_lr}$ scaled by the model’s hidden dimension ($\mathrm{d_{model}}^{-0.5}$) and dynamically adjusts it based on training steps, with a warm-up phase of 500 steps. The learning rate at step $\mathrm{n_{steps}}$ is updated as $\text{lr}=\mathrm{init\_lr}\times\min\left(\mathrm{n_{steps}}^{-0.5},\ \mathrm{n_{warmup}}^{-1.5}\times\mathrm{n_{steps}}\right)$. We set $\beta_{1}=0.9$, $\beta_{2}=0.99$, $\epsilon=1e-8$, batch size 2, and trained for 1 epoch. Additionally, we only train on a randomly sampled subset of 400,000 ECG instances for each respective dataset due to computational resources, unless specified otherwise. We also utilize LoRA (Hu et al., 2021) to finetune the LLM with rank = 16, $\alpha_{LoRA}=32$, and dropout = 0.05. We conduct our experiments on 4 NVIDIA RTX A6000 48 GB GPUS.
During inference, we evaluate our model with number of merges num_merges = 3500, sequence length $L=1024$, and ECG length $T=500$ unless specified otherwise. We use popular metrics for NLG namely the BLEU-4 (Papineni et al., 2002), Rouge-L (Lin, 2004), Meteor (Banerjee and Lavie, 2005), and BertScore F1 (Zhang et al., 2020) metrics.
[…]
tab:main_results Method Trained Dataset Inferenced Dataset BLEU-4 Rouge-L Meteor BertScore F1 ECG-Chat (Zhao et al., 2024) MIMIC-IV ECG Pretrain PTB-XL 11.19 29.93 35.10 - $L_{CL}$ 8.10 ± 0.25 31.36 ± 0.31 27.55 ± 0.36 89.35 ± 0.04 $L_{MIM}$ 6.21 ± 0.22 30.63 ± 0.13 24.91 ± 0.14 90.44 ± 0.04 $L_{MERL}$ (Liu et al., 2024) 10.22 ± 0.25 32.95 ± 0.12 25.60 ± 0.17 89.94 ± 0.01 $L_{CL}$ + $L_{MIM}$ 9.33 ± 0.22 30.45 ± 0.21 24.37 ± 0.36 90.29 ± 0.02 ECG-Byte 11.00 ± 0.19 33.41 ± 0.05 24.95 ± 0.09 90.02 ± 0.01 $L_{CL}$ ECG-QA MIMIC-IV ECG-QA MIMIC-IV 10.22 ± 0.06 38.41 ± 0.48 24.66 ± 0.23 90.42 ± 0.09 $L_{MIM}$ 7.90 ± 0.23 29.28 ± 0.38 19.03 ± 0.11 67.91 ± 0.17 $L_{MERL}$ (Liu et al., 2024) 10.95 ± 0.24 38.18 ± 0.58 26.24 ± 0.36 90.80 ± 0.06 $L_{CL}+L_{MIM}$ 8.57 ± 0.14 34.00 ± 0.25 25.22 ± 0.30 87.72 ± 0.04 ECG-Byte 11.23 ± 0.12 42.49 ± 0.53 27.08 ± 0.15 91.30 ± 0.04 $L_{CL}$ ECG-QA PTB-XL ECG-QA PTB-XL 8.89 ± 0.25 28.63 ± 0.47 18.45 ± 0.31 72.63 ± 0.40 $L_{MIM}$ 15.14 ± 0.28 46.71 ± 0.41 29.64 ± 0.30 92.12 ± 0.10 $L_{MERL}$ (Liu et al., 2024) 13.84 ± 0.19 40.14 ± 0.39 26.24 ± 0.35 91.88 ± 0.09 $L_{CL}+L_{MIM}$ 14.72 ± 0.27 42.88 ± 0.13 28.25 ± 0.27 89.40 ± 0.01 ECG-Byte 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
@@ 5.1 Natural Language Generation
We present our main results in Table LABEL:tab:main_results, comparing ECG-Byte with prior works and self-implemented two-stage pretraining methods. Notably, Zhao et al. (2024) is not directly comparable due to differing data splits and pretraining datasets, though reported metrics use the same datasets. Zhao et al. (2024) train on the full MIMIC-IV ECG Pretrain dataset, finetune on an instruction-tuning dataset for ECG-related conversations, and evaluate on PTB-XL (Wagner et al., 2020) using a unified question: “Could you please help me explain my ECG?” To establish baselines, we implement generic two-stage pretraining methods: $L_{CL}$, $L_{MIM}$, and $L_{CL}+L_{MIM}$. Here, $L_{CL}$ employs contrastive learning (Liu et al., 2024; Gopal et al., 2021; Pham et al., 2024; Kiyasseh et al., 2021), $L_{MIM}$ uses Masked Image Modeling (MIM) (Choi et al., 2023; Na et al., 2024; Yang et al., 2022), and $L_{CL}+L_{MIM}$ combines both (Oh et al., 2022; McKeen et al., 2024). These approaches utilize pretrained CLIP (Radford et al., 2021) and ViT (Dosovitskiy et al., 2021), where ECG signals are transformed into three-channel images for finetuning. We adapt Liu et al. (2024)’s state-of-the-art contrastive method ($L_{MERL}$) for fair comparison. Their most effective model uses a 1D ResNet backbone (He et al., 2015); hence, we employ ResNet101 for direct ECG signal processing. For $L_{CL}$, $L_{MIM}$, $L_{CL}+L_{MIM}$, and $L_{MERL}$, training is conducted on the full, preprocessed MIMIC-IV ECG dataset with a batch size of 64 during the first stage. Implementation details for both training stages are in Appendix B. Table LABEL:tab:main_results demonstrates ECG-Byte’s effectiveness, showing competitive or superior performance across all metrics and datasets compared to other methods. Qualitative examples are provided in Appendix C.2.
5.2 Cross Dataset Transferability
We present the results of cross-dataset transferability in Table LABEL:tab:cross, comparing our approach, ECG-Byte, with two-stage pretraining methods. ECG-Byte achieves the best zero-shot transfer performance from the ECG-QA PTB-XL dataset to the ECG-QA MIMIC-IV dataset. When transferring from the ECG-QA MIMIC-IV dataset to the ECG-QA PTB-XL dataset, although other 2-stage pretraining methods demonstrate higher performance, ECG-Byte maintains competitive results across all metrics.
@@ 5.3 Efficiency of ECG-Byte
We compare the efficiency of our end-to-end approach using ECG-Byte with 2-stage pretraining methods in Table LABEL:tab:efficiency. First, we examine the amount of data required for each method. As previously noted, the first stage of the two-stage pretraining methods is trained on the full MIMIC-IV ECG dataset (Johnson et al., 2023) using segmented ECGs, which are also used as input during the second stage. While ECG-Byte is trained on unsegmented ECGs, we convert the number of unsegmented ECGs to an equivalent count of segmented ECGs. Additionally, the reduced data requirement for ECG-Byte is due to our sampling approach, where only a subset of ECGs is used to train the tokenizer. Under these settings, our proposed method achieves competitive results using approximately $\sim$48of the data required for 2-stage pretraining methods. In terms of training time, our approach requires less than half the time needed for two-stage pretraining. The training time for the two-stage methods is averaged across our self-implemented approaches ($L_{CL}$, $L_{MIM}$, and $L_{CL}+L_{MIM}$) and the $L_{MERL}$ method proposed by Liu et al. (2024).

### v2

@@ 3.2 ECG as Bytes
After training ECG-Byte, we encode any quantized ECG signal $X_{\text{symb}}$ by first converting each byte to a 32-bit unsigned integer and building a trie structure, where each node represents a byte or a merged token sequence from prior encoding steps. The trie is initialized with single-byte tokens (0-255) and is extended with custom token sequences from the learned merge history. For each byte sequence in the input, the encoding function traverses the trie to find the longest match, replacing matched sequences with their assigned token IDs. If no match is found, the byte is added to the output as-is. The final encoded sequence is returned as output_ids, where we will denote as $X_{\text{ID}}$. We provide the detailed pseudocode of the encoding process in Algorithm 4 in the Appendix due to page limitations.
@@ 3.3 Large Language Model
In this study, we utilize the Llama-3.2-1B (Grattafiori et al., 2024) checkpoint through the HuggingFace API (Wolf et al., 2020) unless specified otherwise. We also provide an ablation study in subsection 5.4 where we utilize other popular LLMs, such as GPT2 XL 1.5B (Radford et al., 2019), Gemma 2B (Team et al., 2024), and OPT 1.3B (Zhang et al., 2022). With the exception of modifying the token embedding size (since we add new ECG tokens), we utilize all default hyperparameters out-of-the-box unless specified otherwise.
@@ 4.1 Experimental Settings
We finetuned the LLM using the AdamW optimizer (Kingma and Ba, 2017) with a learning rate of $41e-$ 4, weight decay of $21e-$ 2 (Krogh and Hertz, 1991), and a custom learning rate scheduler. We provide details about the learning rate scheduler in Appendix A.1 due to page limitations. We train the LLMs for 1 epoch and with a batch size of 2. We set the exponential decay rates to be $\beta_{1}=0.9$ and $\beta_{2}=0.99$. We set $\epsilon=1e-8$ as a constant for numerical stability. Due to computational constraints, we limit each dataset to a randomly sampled training subset of 400,000 ECG instances and an independent inference subset of 25,000 instances, unless specified otherwise. We also utilize LoRA (Hu et al., 2021) to finetune the LLM with rank = 16, $\alpha_{LoRA}=32$, and dropout = 0.05. We conduct our experiments on 4 NVIDIA RTX A6000 48 GB GPUs.
During inference, we evaluate our model with number of merges num_merges = 3500, sequence length $L=1024$, and ECG length $T=500$ unless specified otherwise. We use popular metrics for NLG namely the BLEU-4 (Papineni et al., 2002), Rouge-L (Lin, 2004), Meteor (Banerjee and Lavie, 2005), and BertScore F1 (Zhang et al., 2020) metrics. All reported NLG results are presented as means with standard deviations computed over 5 random seeds.
@@
Method Trained Dataset Inferenced Dataset BLEU-4 Rouge-L Meteor BertScore F1
ECG-Chat (Zhao et al., 2024) MIMIC-IV ECG Pretrain PTB-XL 11.19 29.93 35.10 -
$L_{CL}$ 8.10 ± 0.25 31.36 ± 0.31 27.55 ± 0.36 89.35 ± 0.04
$L_{MIM}$ 6.21 ± 0.22 30.63 ± 0.13 24.91 ± 0.14 90.44 ± 0.04
$L_{MERL}$ (Liu et al., 2024a) 10.22 ± 0.25 32.95 ± 0.12 25.60 ± 0.17 89.94 ± 0.01
$L_{Dual}$ 9.33 ± 0.22 30.45 ± 0.21 24.37 ± 0.36 90.29 ± 0.02
ECG-Byte 11.00 ± 0.19 33.41 ± 0.05 24.95 ± 0.09 90.02 ± 0.01
$L_{CL}$ ECG-QA MIMIC-IV ECG-QA MIMIC-IV 10.22 ± 0.06 38.41 ± 0.48 24.66 ± 0.23 90.42 ± 0.09
$L_{MIM}$ 7.90 ± 0.23 29.28 ± 0.38 19.03 ± 0.11 67.91 ± 0.17
$L_{MERL}$ (Liu et al., 2024a) 10.95 ± 0.24 38.18 ± 0.58 26.24 ± 0.36 90.80 ± 0.06
$L_{Dual}$ 8.57 ± 0.14 34.00 ± 0.25 25.22 ± 0.30 87.72 ± 0.04
ECG-Byte 11.23 ± 0.12 42.49 ± 0.53 27.08 ± 0.15 91.30 ± 0.04
$L_{CL}$ ECG-QA PTB-XL ECG-QA PTB-XL 8.89 ± 0.25 28.63 ± 0.47 18.45 ± 0.31 72.63 ± 0.40
$L_{MIM}$ 15.14 ± 0.28 46.71 ± 0.41 29.64 ± 0.30 92.12 ± 0.10
$L_{MERL}$ (Liu et al., 2024a) 13.84 ± 0.19 40.14 ± 0.39 26.24 ± 0.35 91.88 ± 0.09
$L_{Dual}$ 14.72 ± 0.27 42.88 ± 0.13 28.25 ± 0.27 89.40 ± 0.01
ECG-Byte 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
@@ 5.1 Natural Language Generation
We present our main results in Table 1, comparing ECG-Byte with prior works and self-implemented two-stage pretraining methods. Notably, Zhao et al. (2024) is not directly comparable due to differing data splits and pretraining datasets. Zhao et al. (2024) train on the full MIMIC-IV ECG Pretrain dataset, finetune on an instruction-tuning dataset for ECG-related conversations, and evaluate on PTB-XL (Wagner et al., 2020) using a unified question: “Could you please help me explain my ECG?” We also note that ECG-Chat (Zhao et al., 2024) employs numerous generation enhancements such as Retrieval Augmented Generation (RAG), DSPy for automatic prompt tuning, and an adapted Diagnosis-Driven Prompt (DDP) inspired by Jin et al. (2024). In our work, we only observe performance of NLG given a text prompt and ECG signal. Therefore, to establish comparable baselines, we implement generic two-stage pretraining methods utilizing strong encoders: $L_{CL}$, $L_{MIM}$, and $L_{Dual}=L_{CL}+L_{MIM}$. Here, $L_{CL}$ employs contrastive learning (Liu et al., 2024a; Gopal et al., 2021; Pham et al., 2024; Kiyasseh et al., 2021), $L_{MIM}$ uses Masked Image Modeling (MIM) (Choi et al., 2023; Na et al., 2024; Yang et al., 2022), and $L_{Dual}$ combines both (Oh et al., 2022; McKeen et al., 2024). All three implementations utilize pretrained CLIP (Radford et al., 2021) and ViT (Dosovitskiy et al., 2021) models, where ECG signals are transformed into three-channel images for finetuning. Additionally, we adapt Liu et al. (2024a)’s state-of-the-art contrastive method ($L_{MERL}$) for fair comparison. Their most effective model uses a 1D ResNet backbone (He et al., 2015), therefore we utilize the ResNet101 variant. For $L_{CL}$, $L_{MIM}$, $L_{Dual}$, and $L_{MERL}$, training is conducted on the full, preprocessed MIMIC-IV ECG dataset with a batch size of 64 during the first stage to simulate prior work settings. Implementation details for both training stages are in Appendix C. Table 1 demonstrates ECG-Byte’s effectiveness, showing competitive or superior performances across all metrics and datasets compared to other methods. We also highlight that ECG-Byte was initially trained on a separate, unsegmented subset of 200,000 ECGs from the MIMIC-IV ECG dataset. With this fact, ECG-Byte displays strong performances when training and inferencing the LLM on ECGs of different lengths (i.e., segmented to 2 seconds) and from another dataset (i.e., PTB-XL). Qualitative examples of generated answers are provided in Appendix D.4.
@@ 5.2 Efficiency of ECG-Byte
We compare the efficiency of our end-to-end (end-to-end A/B) approach (ECG-Byte A/B) with two-stage pretraining (1st Stage / 2nd Stage) in Table 2. The top and middle sections report data and time requirements for both settings used in our study. The two-stage methods first train on the full MIMIC-IV ECG dataset (Johnson et al., 2023) using segmented ECGs, which are then used as inputs in the second stage. Although ECG-Byte is trained on unsegmented ECGs, we convert these counts to segmented-equivalents and reduce data requirements by sampling only a subset for tokenizer training. Training time for two-stage methods is averaged over our self-implemented approaches ($L_{CL}$, $L_{MIM}$, $L_{Dual}$) and the $L_{MERL}$ method (Liu et al., 2024a). Under these settings, our method achieves competitive results using approximately 48% of the data and is about 3 times faster in total training time. Even when training ECG-Byte on 500,000 unsegmented ECGs (equivalent to 2,500,000 segmented ECGs) as seen in the bottom section (ECG-Byte B), our method remains 2.38 times faster than two-stage pretraining, highlighting its efficiency.
Table 2: Efficiency of our method compared against two-stage pretraining methods.
Method # of Data Total # of Data Time (min.) Total Time (min.)
1st Stage 2,513,435 2,913,435 $\sim$1258.50 $\sim$1727.75
2nd Stage 400,000 $\sim$469.25
ECG-Byte A 1,000,000 1,400,000 $\sim$152.47 $\sim$572.79
end-to-end A 400,000 $\sim$420.32
ECG-Byte B 2,500,000 2,900,000 $\sim$304.27 $\sim$724.59
end-to-end B 400,000 $\sim$420.32
Table 3: Mean results with standard deviations over 5 random seeds on training two-stage pretraining methods in an end-to-end fashion.
Method Trained Dataset Inferenced Dataset BLEU-4 Rouge-L Meteor BertScore F1
$L_{MIM}$ ECG-QA MIMIC-IV ECG-QA MIMIC-IV 7.90 ± 0.23 29.28 ± 0.38 19.03 ± 0.11 67.91 ± 0.17
End-to-End $L_{MIM}$ 7.71 ± 0.36 28.19 ± 0.13 18.88 ± 0.54 67.82 ± 0.04
$L_{MERL}$ (Liu et al., 2024a) 10.95 ± 0.24 38.18 ± 0.58 26.24 ± 0.36 90.80 ± 0.06
End-to-End $L_{MERL}$ (Liu et al., 2024a) 8.26 ± 0.17 34.88 ± 0.26 22.14 ± 0.33 86.64 ± 0.03
ECG-Byte 11.23 ± 0.12 42.49 ± 0.53 27.08 ± 0.15 91.30 ± 0.04
$L_{MIM}$ ECG-QA PTB-XL ECG-QA PTB-XL 15.14 ± 0.28 46.71 ± 0.41 29.64 ± 0.30 92.12 ± 0.10
End-to-End $L_{MIM}$ 12.04 ± 0.17 40.44 ± 0.25 24.84 ± 0.43 87.71 ± 0.11
$L_{MERL}$ (Liu et al., 2024a) 13.84 ± 0.19 40.14 ± 0.39 26.24 ± 0.35 91.88 ± 0.09
End-to-End $L_{MERL}$ (Liu et al., 2024a) 11.27 ± 0.34 38.83 ± 0.17 24.03 ± 0.21 85.64 ± 0.32
ECG-Byte 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
5.3 Training Two-Stage Pretraining Methods in an End-to-End Fashion
Although training methods are conventionally categorized as either two-stage or end-to-end, some works adopt an end-to-end strategy using an ECG-specific encoder architecture typically seen in two-stage methods (Wan et al., 2024). In Table 3, we simulate the approach of Wan et al. (2024) by jointly training the ECG encoder, learnable projection matrix, and LLM end-to-end using only an autoregressive objective. This contrasts with the two-stage method, where the ECG encoder is first trained with a self-supervised learning (SSL) objective and then frozen. We select strong ECG encoders from Table 1—namely, $L_{MIM}$ and $L_{MERL}$—and report performance on the ECG-QA MIMIC-IV and PTB-XL datasets (Oh et al., 2023). We denote the traditional 2-stage training approach for both ECG encoders as $L_{MIM}$ and $L_{MERL}$ and the end-to-end adaptation of 2-stage training approaches as End-to-End $L_{MIM}$ and End-to-End $L_{MERL}$ respectively in Table 3. The results clearly show a performance drop when the ECG encoder is not pretrained on a large dataset. This finding aligns with the Vision-Language Model (VLM) domain, where vision encoders are initially trained on internet-scale image data before integration with an LLM for visual understanding (Liu et al., 2023). Similarly, these findings indicate that ECG signals require extensive first-stage training with a dedicated self-supervised learning objective to learn robust representations before being applied to NLG. Even powerful, general-purpose pretrained encoders like ViT (Dosovitskiy et al., 2021) necessitate an initial training phase to effectively capture the unique characteristics of ECG data. In contrast, our method, ECG-Byte, compresses ECG signals using a rule-based algorithm (e.g., BPE) to represent them as tokens, enabling direct LLM training. Although these tokens belong to a different modality, we believe that the similarity in input representation allows the LLM to adapt more effectively without extensive first-stage training.
5.4 Ablation Study
We conduct several ablation studies to show the variability of performance with ECG-Byte when we use varying LLMs during finetuning, sequence lengths $L$, and ECG lengths $T$. With the exception of the ablating parameter, we fix all other parameters to Llama 3.2 1B, num_merges = 3500, $L=1024$, and $T=500$. We report results when training and testing on the PTB-XL variant of ECG-QA (Oh et al., 2023) unless specified otherwise.
Table 4: Ablation study on using different LLMs.
LLM BLEU-4 Rouge-L Meteor BertScore F1
GPT2 XL 1.5B (Radford et al., 2019) 12.30 ± 0.19 41.33 ± 0.57 26.48 ± 0.33 92.00 ± 0.06
Gemma 2B (Team et al., 2024) 13.78 ± 0.18 45.48 ± 0.55 28.32 ± 0.23 92.01 ± 0.02
OPT 1.3B (Zhang et al., 2022) 12.26 ± 0.20 41.84 ± 0.52 26.21 ± 0.29 91.78 ± 0.04
Llama 3.2 1B (Grattafiori et al., 2024) 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
Different LLMs
We show the variability in performance of ECG-Byte when using different LLMs with similar numbers of parameters in Table 4. While Llama 3.2 1B (Grattafiori et al., 2024) achieves the best results, GPT2 XL 1.5B (Radford et al., 2019), Gemma 2B (Team et al., 2024), and OPT 1.3B (Zhang et al., 2022) also deliver comparable performances. These findings demonstrate that our method is not limited to Llama 3.2 1B but can achieve similar results across a variety of LLMs.
Table 5: Ablation study on varying sequence lengths $L$.
$L$ BLEU-4 Rouge-L Meteor BertScore F1
512 13.61 ± 0.15 48.15 ± 0.57 29.10 ± 0.28 92.41 ± 0.05
1024 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
2048 13.88 ± 0.22 45.21 ± 0.48 28.31 ± 0.27 90.88 ± 0.02
Table 6: Ablation study on varying lengths $T$.
$T$ BLEU-4 Rouge-L Meteor BertScore F1
250 12.64 ± 0.20 47.31 ± 0.26 27.97 ± 0.21 92.32 ± 0.06
500 13.93 ± 0.21 47.08 ± 0.56 29.17 ± 0.31 92.53 ± 0.07
1250 11.01 ± 0.19 43.84 ± 0.28 25.49 ± 0.20 93.07 ± 0.03
2500 14.54 ± 0.17 48.03 ± 0.27 32.11 ± 0.22 92.91 ± 0.04
Sequence Length
Input lengths for LLMs are critical for efficient training since the computation of attention scales quadratically with sequence length (Vaswani et al., 2023). Table 6 presents results for various sequence lengths $L$, which reveal minimal performance differences. As Figure 2 shows, most ECGs are encoded with token lengths around 500. Thus, we hypothesize that even small $L$ values (e.g., $L=512$) preserve the complete ECG information, resulting in only slight performance variations. We note that this can change depending on the initial ECG length $T$, in which we ablate in the next paragraph. We leave it up to future works for finding the sweet spot for the sequence length $L$ and its interaction with the initial ECG length $T$.